OpenStudy (help!!!!):

8. If Sam can do a job in 4 days that Lisa can do in 6 days and Tom can do in 2 days, how long would the job take if Sam, Lisa, and Tom worked together to complete it? A. 0.8 days B. 1.09 days C. 1.23 days D. 1.65 days E. 1.97 days

OpenStudy (help!!!!):

i know you first do 1/4+1/6+1/2=11/12

OpenStudy (raden):

4*6*2 / (4*6 + 4*2 + 6*2) = ...

OpenStudy (help!!!!):

is there another way to do it? For example you add 1/4+2/6+1/2=11/12 then..........

OpenStudy (help!!!!):

is there another way to do this problem?

OpenStudy (anonymous):

Solve the following for x:\[\frac{1}{4}+\frac{1}{6}+\frac{1}{2}= \frac{1}{x}\]

OpenStudy (help!!!!):

thank you so much!!

OpenStudy (anonymous):

You're welcome.